OpenStudy (anonymous):

how do I solve triangle x^2 +6^2=10^2 for x

OpenStudy (anonymous):

ok so we will get the following... x^2 = (10^2) - (6^2) x^2 = 100 - 36 x^2 = 64 so x = 8